24/+3Cl₂ → 2AlCl3 In an experiment, 54 grams of aluminium reacted with excess chlorine gas. How many formula units of AICI, are produced?


1
Expert's answer
2022-03-07T10:32:03-0500

Solution:

Calculate the moles of aluminium (Al):

The molar mass of Al is 26.98 g/mol

Therefore,

Moles of Al = (54 g Al) × (1 mol Al / 26.98 g Al) = 2.00 mol Al


Balanced chemical equation:

2Al + 3Cl2 → 2AlCl3

According to stoichiometry:

2 mol of Al produces 2 mol of AlCl3

Thus, 2.00 mol of Al produces:

(2.00 mol Al) × (2 mol AlCl3 / 2 mol Al) = 2.00 mol AlCl3


One mole of any substance contains 6.022×1023 formula units.

Therefore,

Formula units of AICI3 = (2.00 mol AlCl3) × ( 6.022×1023 formula units / 1 mol) = 1.2×1024 formula units


Answer: 1.20×1024 formula units of AICI3 are produced